Polonium-214 has a short half-life of 164 seconds. How many seconds would it take for 8.0g of this isotope to decay to 0.25 g?
1
Expert's answer
2013-03-13T08:53:45-0400
You can answer this using half-life time:

t1/2 = 164 seconds

t1/2 = ln2/k

k is rate of the reaction
164 = 0.693 / k
k = 0.00423 s-1
ln 0.25 / 8.0 = -0.00423 t
-3.47 = -0.00423 t
t = 820 s